What is the area under the curve to the right of z= -1.35? Provide solution and drawing.
1
Expert's answer
2021-09-24T02:56:03-0400

Z= - 1.35, Asea to the sightof Z=P(Z>Z)

P(Z>-1.35)= 1 - P(Z ≤ -1.35)= 1 - 0.0005 (By Z table)

= 0.9115